Find the most economical dimensions for a closed cylindrical can containing a quart.


1
Expert's answer
2021-01-14T15:50:53-0500

It sounds like "most economical dimensions" means you want to use the least amount of material for the can so that it still contains a quart. In other words, you're minimizing the surface area of the can.


Here is not given the volume "V" of cylinder is "V=\pi \ r^2 \ h=\frac{1}{4} \ \pi \ d^2 \ h=1 quart" . I applied differentiation like this:


"\frac{1}{4}\pi [d^2\frac{dh}{dd}+2dh]=0;"


"\frac{dh}{dd}=-\frac{2h}{d};"


Total area (closed both ends):


"A_T=2(\frac{1}{4}\pi d^2)+\pi dh" ;


"A_T=\frac{1}{2}\pi d^2+\pi dh;"


"\frac{dA_T}{dd}=πd+π[d\frac{dh}{dd}+h]=0;"

"dd\frac{dh}{dd}+h=0;"


"d+d(\frac{−2h}{d})+h=0;"


"d=h;"


Result: The diameter of the base of the can is equal to the height of this can.